OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form. Answer 3x + y = 9 3x – 3y = –8 x – 3y = –8 x – 3y = –9

OpenStudy (amistre64):

dont delete the questions. thats considered to be tabboo behaviour by the people who built the site.

OpenStudy (anonymous):

Tabboo and im new to this site lol sorry

OpenStudy (amistre64):

:) just keep in mind that the sites motivation is for the community to benefit from the interactions; not just the person asking the questions.

OpenStudy (amistre64):

if we subtract the point we can establish a slope to judge by

OpenStudy (amistre64):

(4, 4) (1, 3) ----- 3 1 ; slope = 1/3

OpenStudy (anonymous):

thats how u do it holy cow how easy

OpenStudy (amistre64):

the bottom 3 tells me that y has a 3

OpenStudy (amistre64):

lets test em x – 3y = –8 1 3 1 - 9 = -8 .. yes x – 3y = –9 1 3 1 - 9 = -9 ...no go with the 3rd